# CoDefeater: Using LLMs To Find Defeaters in Assurance Cases

### v1

@@ 1. Introduction
Safety-critical systems have become deeply integrated into many societal domains, including healthcare, transportation, energy, and aviation (47; 22; 33). Failures in these systems can lead to catastrophic consequences for human safety, including fatalities and environmental and property damage (60). This has led to an increased focus on their dependence, reliability, and safety (58; 39). Many systems must comply with regulations (55; 10), provide evidence of safety, and undergo rigorous certification processes (36; 54) for approval from regulatory bodies. Assurance cases (ACs) have emerged as a common practice for this purpose, facilitating the verification of system correctness and the validation of specific claims regarding safety, security, and trustworthiness, among others (46; 8; 6).
An assurance case is a structured hierarchy of claims and arguments supported by evidence that a system will function as intended in a specified environment (61; 8). Several formal notations (e.g., Goal Structuring Notation (GSN) (37), Claims-Arguments-Evidence (CAE) (2), and Eliminative Argumentation (EA) (29)), along with tools (21; 65; 46), have been proposed. However, concerns arise over their completeness, uncertainty, and soundness for cyber-physical systems. (31; 25), leading to false confidence and catastrophic failures (64). For example, failure of the minimum safe altitude warning system that led to a major aviation accident was attributed to incomplete and flawed reasoning in the safety case (32).
@@
Various approaches have been proposed to identify and mitigate defeaters (30; 38). However, manually creating defeaters is a labor-intensive and time-consuming process (52; 49), relying heavily on safety analysts’ judgment, experience, creativity, and understanding of the system. This can lead to confirmation bias (65; 7). As assurance cases evolve with new standards and technological advances, ongoing efforts are focused on formal and semi-automated approaches for detecting and managing defeaters (23; 41). Best practices build assurance cases incrementally, so automating all/or part of this process is important (68; 23).
Large Language Models (LLMs) are increasingly automating software engineering tasks like test generation and defect detection (35; 3). In particular, these models have become valuable in tasks requiring complex understanding, such as vulnerability detection, requirements elicitation, and code generation (69; 17; 5). Moreover, LLMs have shown promise in automating evaluation tasks and acting as surrogate evaluators (18; 70). Consequently, we explore whether LLMs’ capabilities can be harnessed to automate defeater analysis towards the completeness, soundness, and confidence of assurance cases. Despite calls for more research into LLMs’ ability to identify defeaters (62; 66), no study has been conducted to evaluate and investigate their effectiveness.
@@ 2. Background and Related Work
Assurance case arguments typically adopt an inductive approach, where sub-claims offer direct evidence to support the parent claim but do not ensure it with certainty (8; 13). Focusing solely on proving a claim may introduce confirmation bias, as exemplified by the Nimrod aircraft crash (12). Recent approaches have therefore embraced defeasible reasoning, which acknowledges that arguments about system properties in practice are inherently defeasible (28; 43; 57). Defeaters are potential doubts or objections that challenge the validity of a claim, reflecting gaps in evidence and reasoning (28; 44; 25; 51). Figure 2 shows a fragment of an assurance case for an sUAS battery, with examples of defeaters (red boxes).
@@
In software engineering, LLMs are increasingly employed to assist developers and automate tasks such as discovering requirements, code generation, testing, and program synthesis (35; 5). Diemert and Weber (24) have reported ChatGPT’s effectiveness in hazard analysis for safety-critical systems, highlighting their potential to assist human analysts. In the context of software assurance cases, Sivakumar et al. (63) assessed LLM’s (GPT-4) proficiency in understanding GSN representations and its performance in constructing safety cases. Viger et al. (66) proposed using LLMs to identify defeaters; however, they did not report an empirical evaluation of their capabilities. Shahandashti et al. (62) explored LLM’s (GPT-4) understanding of EA notation and defeater concepts, with empirical validation left as future work. We aim to address this gap by examining LLM performance in identifying defeaters for assurance cases in two case studies.
@@ 3.1. Experimental Setup
Datasets. We performed our experiments on two assurance cases. The first is for the CERN Large Hadron Collider (LHC) Machine Protection System (MPS), which provides assurance that the MPS will prevent damage to the LHC from unstable, high-energy particle beams (59; 50). The assurance case for the CERN LHC uses EA notation, and we extracted claim nodes with their corresponding defeaters for our experiments. The second assurance case is a fragment of a larger one that our team has recently created for small Uncrewed Aircraft Systems (sUAS)111Anonymized for submission using GSN. It addresses the claim “the sUAS has enough charge in its battery to complete the mission” (see claim 1.2 in Figure 2). We include both the assurance case fragment and corresponding defeaters in the supplementary material222https://gitlab.com/anonymousdot/codefeater. The real-world complexity of these systems and the availability of defeaters made them suitable for our preliminary experiments. Table 1 provides a summary of the assurance cases.
@@ 3.3. Evaluation Criteria
Due to the complexity and open-ended nature of the task (e.g., varying response length vs. ground truth, and subjectivity), automatic evaluation metrics were not suitable (18). Therefore, we relied on human evaluation for assessing LLM performance. For RQ1, we used a deductive coding approach (45; 24), where the first two authors independently reviewed responses and categorized them as complete match, partial match, or no match based on similarity to ground-truth defeaters. We used the defeaters in the LHC assurance case as the ground truth. For the battery assurance case, a set of ground-truth defeaters was provided by one of the authors (independently) familiar with the domain, following best practices (50). For RQ2, the responses were evaluated for being reasonable (14), i.e., the defeater could reasonably be in the ground truth but had been overlooked. This aimed to assess the LLM’s capability to identify novel defeaters. Figure 4 shows an example of each type of match. Next, the reviewers met to discuss and finalize their assigned codes. In the case of post-discussion disagreement, if one reviewer labeled a response as a partial match and the other as a complete match, we categorized it as a partial match to avoid confirmation bias (26). In the one instance where one reviewer indicated no match while the other identified a partial or complete match, it was discarded. Last, we calculated inter-rater agreement using Cohen’s Kappa (19) to evaluate the consistency and reliability of the coding process.
@@ 4. Results
Figure 5 presents the results of our closed coding process, which assessed the GPT-3.5 model’s performance in identifying defeaters for the two real-world assurance cases in a zero-shot setting. Our experiments showed that the model demonstrated unexpectedly good zero-shot capabilities for defeater analysis. Specifically, it completely identified more than half of all defeaters and partially identified more than a third in both datasets. Fewer than of the defeaters created by human analysts were not identified at all by the LLM. These results were noteworthy given the complexity of the systems under study and the lack of system information, domain knowledge, or few-shot examples provided to the LLM. Additionally, the total coded responses achieved high inter-rater agreement, with a Cohen’s Kappa score of 0.81, (19; 48). Finally, all of these defeaters were identified in the first prompt, illustrating the convenience of the approach. Among the unidentified defeaters, some required specific domain or system knowledge not provided to the model, indicating areas for potential improvement.
@@
We conducted a manual analysis of the unidentified defeaters (n = 17) to investigate whether there were any patterns behind the LLM’s failure. Interestingly, we found that the model struggled to identify those defeaters that implicitly challenged the truth of an assumption. For example, for the claim ”The BICs will not produce a FALSE BEAM-PERMIT to trigger a beam dump, unless a loss of the high-frequency signal (10Mhz) in either Beam Permit Loop (A and B) is detected...”, a (ground-truth) defeater in the LHC assurance case questions the assumption that the 10Mhz signal is the right signal to monitor. In other words, the presence of another similar high-frequency signal might lead to a false indication of TRUE BEAM-PERMIT. Unlike the analysts, the LLM did not question the underlying assumption in the claim and thus did not identify the defeater. Consequently, future work should explore integrating external knowledge sources since, in similar tasks, it significantly enhances LLM’s performance (72; 56).
@@ 5. Discussion
Designing better prompts. Prompt designing has been shown to significantly impact LLM performance (42; 15). Many prompting methods have been proposed, and further investigation for suitability to defeater analysis is needed. Moreover, our study revealed that the LLM can generate creative, redundant, and far-fetched scenarios (e.g., defeaters due to budget constraints). Balancing LLM creativity with defeater relevance poses an important challenge.
@@ Acknowledgements.
- Shahandashti et al. (2024) Kimya Khakzad Shahandashti, Mithila Sivakumar, Mohammad Mahdi Mohajer, Alvine B Belle, Song Wang, and Timothy C Lethbridge. 2024. Evaluating the Effectiveness of GPT-4 Turbo in Creating Defeaters for Assurance Cases. arXiv preprint arXiv:2401.17991 (2024).

### v2

@@ 1. Introduction
Safety-critical systems have become deeply integrated into many societal domains, including healthcare, transportation, energy, and aviation (48; 22; 33). Failures in these systems can lead to catastrophic consequences for human safety, including fatalities and environmental and property damage (61). This has led to an increased focus on their dependence, reliability, and safety (59; 40). Many systems must comply with regulations (56; 10), provide evidence of safety, and undergo rigorous certification processes (36; 55) for approval from regulatory bodies. Assurance cases (ACs) have emerged as a common practice for this purpose, facilitating the verification of system correctness and the validation of specific claims regarding safety, security, and trustworthiness, among others (47; 8; 6).
An assurance case is a structured hierarchy of claims and arguments supported by evidence that a system will function as intended in a specified environment (62; 8). Several formal notations (e.g., Goal Structuring Notation (GSN) (37), Claims-Arguments-Evidence (CAE) (2), and Eliminative Argumentation (EA) (29)), along with tools (21; 65; 47), have been proposed. However, concerns arise over their completeness, uncertainty, and soundness for cyber-physical systems (31; 25), leading to false confidence and catastrophic failures (64). For example, failure of the minimum safe altitude warning system that led to a major aviation accident was attributed to incomplete and flawed reasoning in the safety case (32).
@@
Various approaches have been proposed to identify and mitigate defeaters (30; 38). However, manually creating defeaters is a labor-intensive and time-consuming process (53; 50), relying heavily on safety analysts’ judgment, experience, creativity, and understanding of the system. This can lead to confirmation bias (65; 7). As assurance cases evolve with new standards and technological advances, ongoing efforts are focused on formal and semi-automated approaches for detecting and managing defeaters (23; 42). Best practices build assurance cases incrementally, so automating all/or part of this process is important (68; 23).
Large Language Models (LLMs) are increasingly automating software engineering tasks like test generation and defect detection (35; 3). In particular, these models have become valuable in tasks requiring complex understanding, such as vulnerability detection, requirements elicitation, and code generation (69; 17; 5). Moreover, LLMs have shown promise in automating evaluation tasks and acting as surrogate evaluators (18; 70). Consequently, we explore whether LLMs’ capabilities can be harnessed to automate defeater analysis toward the completeness, soundness, and confidence of assurance cases. Despite calls for more research into LLMs’ ability to identify defeaters (39; 66), no study has been conducted to evaluate and investigate their effectiveness.
@@ 2. Background and Related Work
Assurance case arguments typically adopt an inductive approach, where sub-claims offer direct evidence to support the parent claim but do not ensure it with certainty (8; 13). Focusing solely on proving a claim may introduce confirmation bias, as exemplified by the Nimrod aircraft crash (12). Recent approaches have therefore embraced defeasible reasoning, which acknowledges that arguments about system properties in practice are inherently defeasible (28; 44; 58). Defeaters are potential doubts or objections that challenge the validity of a claim, reflecting gaps in evidence and reasoning (28; 45; 25; 52). Figure 2 shows a fragment of an assurance case for an sUAS battery, with examples of defeaters (red boxes).
@@
In software engineering, LLMs are increasingly employed to assist developers and automate tasks such as discovering requirements, code generation, testing, and program synthesis (35; 5). Diemert and Weber (24) have reported ChatGPT’s effectiveness in hazard analysis for safety-critical systems, highlighting their potential to assist human analysts. In the context of software assurance cases, Sivakumar et al. (63) assessed LLM’s (GPT-4) proficiency in understanding GSN representations and its performance in constructing safety cases. Viger et al. (66) proposed using LLMs to identify defeaters; however, they did not report an empirical evaluation of their capabilities. Khakzad Shahandashti et al. (39) explored LLM’s (GPT-4) understanding of EA notation and defeater concepts, with empirical validation left as future work. We aim to address this gap by examining LLM performance in identifying defeaters for assurance cases in two case studies.
@@ 3.1. Experimental Setup
Datasets. We performed our experiments on two assurance cases. The first is for the CERN Large Hadron Collider (LHC) Machine Protection System (MPS), which provides assurance that the MPS will prevent damage to the LHC from unstable, high-energy particle beams (60; 51). The assurance case for the CERN LHC uses EA notation, and we extracted claim nodes with their corresponding defeaters for our experiments. The second assurance case is a fragment of a larger one that our team has recently created for small Uncrewed Aircraft Systems (sUAS)111Anonymized for submission using GSN. It addresses the claim “the sUAS has enough charge in its battery to complete the mission” (see claim 1.2 in Figure 2). We include both the assurance case fragment and corresponding defeaters in the supplementary material222https://gitlab.com/anonymousdot/codefeater. The real-world complexity of these systems and the availability of defeaters made them suitable for our preliminary experiments. Table 1 provides a summary of the assurance cases.
@@ 3.3. Evaluation Criteria
Due to the complexity and open-ended nature of the task (e.g., varying response length vs. ground truth, and subjectivity), automatic evaluation metrics were not suitable (18). Therefore, we relied on human evaluation for assessing LLM performance. For RQ1, we used a deductive coding approach (46; 24), where the first two authors independently reviewed responses and categorized them as complete match, partial match, or no match based on similarity to ground-truth defeaters. We used the defeaters in the LHC assurance case as the ground truth. For the battery assurance case, a set of ground-truth defeaters was provided by one of the authors (independently) familiar with the domain, following best practices (51). For RQ2, the responses were evaluated for being reasonable (14), i.e., the defeater could reasonably be in the ground truth but had been overlooked. This aimed to assess the LLM’s capability to identify novel defeaters. Figure 4 shows an example of each type of match. Next, the reviewers met to discuss and finalize their assigned codes. In the case of post-discussion disagreement, if one reviewer labeled a response as a partial match and the other as a complete match, we categorized it as a partial match to avoid confirmation bias (26). In the one instance where one reviewer indicated no match while the other identified a partial or complete match, it was discarded. Last, we calculated inter-rater agreement using Cohen’s Kappa (19) to evaluate the consistency and reliability of the coding process.
@@ 4. Results
Figure 5 presents the results of our closed coding process, which assessed the GPT-3.5 model’s performance in identifying defeaters for the two real-world assurance cases in a zero-shot setting. Our experiments showed that the model demonstrated unexpectedly good zero-shot capabilities for defeater analysis. Specifically, it completely identified more than half of all defeaters and partially identified more than a third in both datasets. Fewer than of the defeaters created by human analysts were not identified at all by the LLM. These results were noteworthy given the complexity of the systems under study and the lack of system information, domain knowledge, or few-shot examples provided to the LLM. Additionally, the total coded responses achieved high inter-rater agreement, with a Cohen’s Kappa score of 0.81, (19; 49). Finally, all of these defeaters were identified in the first prompt, illustrating the convenience of the approach. Among the unidentified defeaters, some required specific domain or system knowledge not provided to the model, indicating areas for potential improvement.
@@
We conducted a manual analysis of the unidentified defeaters (n = 17) to investigate whether there were any patterns behind the LLM’s failure. Interestingly, we found that the model struggled to identify those defeaters that implicitly challenged the truth of an assumption. For example, for the claim ”The BICs will not produce a FALSE BEAM-PERMIT to trigger a beam dump, unless a loss of the high-frequency signal (10Mhz) in either Beam Permit Loop (A and B) is detected...”, a (ground-truth) defeater in the LHC assurance case questions the assumption that the 10Mhz signal is the right signal to monitor. In other words, the presence of another similar high-frequency signal might lead to a false indication of TRUE BEAM-PERMIT. Unlike the analysts, the LLM did not question the underlying assumption in the claim and thus did not identify the defeater. Consequently, future work should explore integrating external knowledge sources since, in similar tasks, it significantly enhances LLM’s performance (72; 57).
@@ 5. Discussion
Designing better prompts. Prompt designing has been shown to significantly impact LLM performance (43; 15). Many prompting methods have been proposed, and further investigation for suitability to defeater analysis is needed. Moreover, our study revealed that the LLM can generate creative, redundant, and far-fetched scenarios (e.g., defeaters due to budget constraints). Balancing LLM creativity with defeater relevance poses an important challenge.
@@ Acknowledgements.
- Khakzad Shahandashti et al. (2024) Kimya Khakzad Shahandashti, Mithila Sivakumar, Mohammad Mahdi Mohajer, Alvine Boaye Belle, Song Wang, and Timothy Lethbridge. 2024. Assessing the Impact of GPT-4 Turbo in Generating Defeaters for Assurance Cases. In Proceedings of the 2024 IEEE/ACM First International Conference on AI Foundation Models and Software Engineering (Lisbon, Portugal) (FORGE ’24). Association for Computing Machinery, New York, NY, USA, 52–56. https://doi.org/10.1145/3650105.3652291